A piece of wire is shaped to enclose a square whose area is 169 cm2. It is then reshaped to enclosed a rectangle whose length is 5 cm. how many square centimeters(cm2) is the area of the rectangle?


1
Expert's answer
2021-05-11T13:59:02-0400

length of the square "=\sqrt{169}=13cm"

perimeter "=13\times4=52cm"

for the rectangle, perimeter "=52=2(l+w)"

"52=2(5+w)"

"26=5+w, w=26-5=21cm"

"area=21\times5=105cm^2"